### v5

@@ 1 Introduction
Large language models (LLMs), such as GPT-3, OPT, and LLaMA demonstrate exceptional proficiency in a variety of natural language processing (NLP) tasks and have become key components in applications like chatbots and question-answering systems Brown et al. (2020); Zhang et al. (2022); Chowdhery et al. (2022); Touvron et al. (2023a; b). However, their substantial number of parameters creates significant challenges in deploying these models for real-world services, especially due to the increased memory consumption and computational demands. This limitation restricts their widespread use. Consequently, it is critical to develop techniques that improve the compactness and processing efficiency of LLMs while preserving their linguistic prowess.
Network pruning is a technique aimed at reducing the size and complexity of neural networks by eliminating redundant weight parameters LeCun et al. (1989); Hassibi et al. (1993); Han et al. (2015). Its application in LLMs has been somewhat limited, primarily due to challenges that arise in managing sparse matrices Frantar & Alistarh (2022); Sun et al. (2023); Frantar & Alistarh (2023); Zhang et al. (2024). This complexity becomes particularly evident when using modern GPU hardware, as these systems are typically optimized for operations involving dense matrices Wang (2020); Shi et al. (2020).
@@
In the realm of LLMs, a significant similarity in output is observed among successive transformer blocks Din et al. (2023); Liu et al. (2023). This similarity arises because each transformer block incrementally contributes to the residual path spanning the entire LLM. Figure 1 depicts the typical architecture of conventional LLMs, characterized by a continuous stack of transformer blocks. A key aspect of LLM computation is the residual path that traverses the entire network, a feature introduced to stabilize the backpropagation process during training. Consequently, each transformer block contributes its computational outputs, derived from both the attention mechanisms and feed-forward layers, to this residual path. This design feature results in a considerable degree of output similarity between consecutive transformer blocks, leading to redundancy within the LLMs.
@@ 2.1 Pruning
Standard LLMs are designed as shown in Figure 1, with a series of identically structured transformer blocks arranged sequentially. Each transformer block in these models consists of attention and feed-forward layers. The operations in both the linear and feed-forward layers primarily involve matrix multiplication between the weight parameters and input activations. Given this structure, there are two main strategies for developing compact and fast LLMs: enhancing the efficiency of each individual transformer block or reducing the overall number of blocks. To improve both the compactness and processing speed of individual blocks, pruning is frequently utilized LeCun et al. (1989); Hassibi et al. (1993); Han et al. (2015). This technique involves the strategic removal of superfluous weight parameters, thereby reducing the computational load associated with matrix multiplication in each linear and feed-forward layer. However, the application of pruning techniques to LLMs encounters a notable challenge.
Challenge 1) Limitation in Achieving LLM Inference Speedup: Pruning can be divided into two main types: unstructured and structured Frantar & Alistarh (2022); Sun et al. (2023); Frantar & Alistarh (2023); Zhang et al. (2024). Unstructured pruning targets the removal of individual weights, leading to sparse weight matrices within the model. This sparsity can create complex data access patterns, complicating the management of sparse data and potentially hindering the model’s acceleration. Particularly with NVIDIA GPUs, achieving a speedup through unstructured pruning typically requires reaching a high level of sparsity, often more than 90% Wang (2020); Shi et al. (2020). However, for LLMs, achieving pruning ratios above 50% is often difficult without substantially compromising their linguistic prowess. Structured pruning, on the other hand, involves the elimination of predefined units of weights to create more hardware-friendly patterns, such as 2:4 pruning Sun et al. (2023); Frantar & Alistarh (2023); Zhang et al. (2024) or channel-wise (a.k.a. row/column-wise) pruning Ashkboos et al. (2024); Ma et al. (2023). The goal of structured pruning is to form dense matrices that are more efficiently processed by hardware, thereby enhancing efficiency. Nonetheless, the anticipated speedup from structured pruning, ideally proportional to the pruning ratio, frequently fails to be realized in practical settings.
@@
Recent NVIDIA GPUs support the acceleration of 2:4 fine-grained structured sparsity using their Tensor Cores Mishra et al. (2021). This form of sparsity involves having two zero values within each contiguous block of four values, resulting in a natural sparsity rate of 50%. The peak performance of sparse tensor cores is twice that of dense tensor cores. Thus, previous research on pruning LLMs has extensively examined 2:4 pruning techniques Sun et al. (2023); Frantar & Alistarh (2023); Zhang et al. (2024). However, it is challenging to achieve the desired speedup with 2:4 pruning. The speedup achieved with 2:4 pruning heavily depends on the size of the input matrix. Figure 3 provides a comparative analysis of matrix multiplication latency on an NVIDIA RTX A6000 GPU, using both NVIDIA cuBLAS and CUTLASS libraries. The matrix multiplication involves batched inputs with dimensions of $b\times h$, along with a weight matrix size of $h\times h$, where $b$ represents the batch size and $h$ represents the hidden size. The measurement results indicate that within a realistic batch size range of 1 to 128, 2:4 pruning does not result in a speedup. However, as the batch size is increased beyond this range, 2:4 pruning begins to show a speedup, eventually reaching approximately 1.5$\times$ the speed of dense matrix multiplication. To achieve significant speedups, it is essential to have a sufficiently large matrix size, which can lead to high arithmetic intensity Mishra et al. (2021). In summary, achieving faster LLM inference with conventional pruning methods has two main limitations. Firstly, it requires dedicated hardware, such as the 2:4 sparse matrix multiplication accelerator adapted in recent NVIDIA GPUs, to enhance inference speed. Secondly, even with dedicated hardware support, the extent of speedup achievable through 2:4 pruning is influenced by factors like batch size and model size.
@@
As an alternative to traditional pruning methods, approaches like LLM-Pruner Ma et al. (2023) and SliceGPT Ashkboos et al. (2024) have been developed, focusing on the elimination of redundant neurons and the weight parameters connected to them. These methods involve the elimination of the entire channel (row/column) from weight matrices, allowing pruned weights to remain in a dense format. This retention of dense matrices enables direct speedup in matrix multiplication. However, it is challenging to maintain linguistic capabilities of LLMs with these channel-wise pruning approaches without the help of extensive fine-tuning Ma et al. (2023). Additionally, speedup in matrix multiplication does not always lead to corresponding improvement in the end-to-end inference speed of LLMs Ashkboos et al. (2024). This limitation arises because transformer blocks in LLMs encompass various operations beyond matrix multiplication that involve weight parameters. These operations include, but are not limited to, layer normalization and attention mechanisms (Figure 1).
Recently, Deja Vu Liu et al. (2023) has introduced a new pruning approach termed contextual sparsity. This method dynamically assesses whether to bypass certain segments of a layer operation, based on the context of the input. It has shown effectiveness in accelerating inference, particularly in single-batch scenarios. However, its dynamic nature introduces a set of challenges akin to those encountered with early exit strategies. These challenges will be further examined and discussed in the following subsection.
@@ 2.2 Early Exit
Reducing the total number of transformer blocks in the LLM can directly lead to increased processing speed, proportional to the reduction in the number of blocks. A notable technique in this context is the early exit Schuster et al. (2022); Din et al. (2023); Varshney et al. (2023); Chen et al. (2023). This method dynamically assesses whether to process subsequent transformer blocks. When the model reaches a certain level of confidence in its early outputs, it can halt further processing, effectively reducing computational cost. Another recent approach Del Corro et al. (2023) proposes an alternative method for bypassing transformer blocks. This strategy suggests that removing blocks, particularly those at the beginning of LLMs, might be more feasible. However, all of these strategies, which involve skipping transformer blocks, typically require either dynamic decision-making or extensive training to be effective. To assess the impact of such removals without dynamic decision-making and training, we analyze the perplexity scores of OPT-6.7B/13B on the WikiText-2 dataset Merity et al. (2016) after removing various consecutive blocks, as illustrated in Figure  4. The analysis results show that the “Early Exit” approach leads to a noticeable decline in the linguistic prowess of the LLMs, as evidenced by a significant increase in their perplexity scores. Moreover, “Chunk Best”, which represents the best perplexity results achieved by testing all possible removable points of consecutive blocks, also results in a significant increase in perplexity as the number of removed blocks increases. Therefore, the concept of removing consecutive blocks from LLMs is not effective without dynamic decision-making and training. In addition, applying early exit to LLMs, particularly in real-world implementation, faces the following challenges.
@@ 3.1 Output Similarity across Transformer Blocks
Here, $x_{i}$ denotes the output of the $i$-th transformer block, and $T_{i}$ denotes the operation results of the $i$-th transformer block. To comprehensively analyze and quantify the output similarity within the transformer blocks, we analyze cosine similarity between outputs of these blocks. This analysis is based on the distance measurement method outlined in Din et al. (2023). It involves feeding a single random token into each LLM and measuring the cosine similarity between the output of the $i$-th transformer block and that of the $j$-th transformer block is defined as follows:
@@ 4.1 Experimental Setup
We implement SLEB in PyTorch Paszke et al. (2019), using the HuggingFace Transformers library Wolf et al. (2020). The experiments on redundancy verification and elimination of transformer blocks are executed on NVIDIA A100 GPUs equipped with 80GB of memory. SLEB requires 2 GPUs for pruning OPT-66B and LLaMA-2-70B, and 1 GPU for pruning smaller models. Runtime of SLEB for pruning LLMs falls within the range of runtime observed with other pruning methods (See Appendix A.1). For example, SLEB can fully compress LLaMA-2-70B, within approximately 1.5 hours. SLEB is processed based on the inference-only approach without any fine-tuning, aligning with contemporary practices in post-training pruning of LLMs. We use 128 samples randomly selected from WikiText-2 training dataset as calibration data, following the approach used in previous works Ashkboos et al. (2024).
Our evaluation encompasses models from the OPT and LLaMA-2 families. We assess SLEB under two target sparsity levels: 10% and 20%. In case that the product of the total number of transformer blocks in a model and the target sparsity is not an integer, we round up the product to determine the number of transformer blocks to remove. The proposed SLEB is compared to previous works that are widely recognized for their potential to accelerate LLM inference. These previous works include SparseGPT Frantar & Alistarh (2023), Wanda Sun et al. (2023), and DSnoT Zhang et al. (2024), which utilize 2:4 structured pruning methods, and LLM-Pruner and SliceGPT, which employ channel-wise pruning.
@@ 4.3 Language Modeling
We evaluate the linguistic capabilities of pruned LLMs on language modeling tasks. Table 1 showcases the perplexity results on C4 validation dataset Raffel et al. (2020). It is important to note that though the sparsity of SLEB is 10% to 20% whereas previous works show a sparsity ratio between 20% to 50%, as we will discuss in Section 4.6, SLEB outperforms in LLM inference speedup. Though Wanda and DSnoT completely fail on OPT-66B, SLEB sufficiently preserves perplexity scores for all evaluated cases. Moreover, SLEB outperforms previous pruning methods in the majority of cases by a significant margin, though it uses a coarser granularity of pruning unit. These results once again emphasize a point discussed in Section 3.1 - LLMs exhibit robustness at the level of transformer blocks. By effectively identifying and removing transformer blocks, SLEB can have a minimal impact on perplexity scores. The perplexity results on WikiText-2 test dataset can be found in Appendix B.2.
@@ 4.5 Zero-shot Tasks
Following the evaluation methods presented in previous works Ashkboos et al. (2024), we evaluate SLEB on following zero-shot tasks: PIQA Bisk et al. (2020), WinoGrande Sakaguchi et al. (2019), HellaSwag Zellers et al. (2019), ARC-easy and ARC-challenge Clark et al. (2018). LM Evaluation Harness Gao et al. (2023) with its default parameters is used for the evaluation.
@@ 4.7 Compatibility with Post-Training Quantization
Post-training quantization (PTQ) is a well-established technique for compressing and accelerating LLMs Frantar et al. (2023); Lee et al. (2024); Lin et al. (2023). If PTQ can be applied to compress LLMs alongside SLEB, we can expect further improvements in memory efficiency and LLM inference speed. In Figure 10, we present a comparison of perplexity scores for LLMs that have undergone SLEB with target sparsity of 20% and those that have been further compressed with AWQ Lin et al. (2023), a state-of-the-art PTQ algorithm that quantizes weights to 4-bit integer values. The experimental results indicate that applying PTQ to streamlined models has no discernible impact on their perplexity. More analysis results on compatibility with PTQ are provided in Appendix B.6.
@@ B.8 Fine-tuning
We investigate the effects of fine-tuning on LLaMA-2-7B after pruning with 2:4 pruning (SparseGPT), channel-wise pruning (SliceGPT), and SLEB. All pruned models are trained fore a single epoch on the Alpaca training dataset Taori et al. (2023). For the efficient fine-tuning procedure, we employ LoRA Hu et al. (2021) modules with a rank of 8.

### v6

@@ 1 Introduction
Large language models (LLMs), such as GPT-3, OPT, and LLaMA demonstrate exceptional proficiency in a variety of natural language processing (NLP) tasks and have become key components in applications like chatbots and question-answering systems (Brown et al., 2020; Zhang et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023a, b). However, their substantial number of parameters creates significant challenges in deploying these models for real-world services, especially due to the increased memory consumption and computational demands. This limitation restricts their widespread use. Consequently, it is critical to develop techniques that improve the compactness and processing efficiency of LLMs while preserving their linguistic prowess.
Network pruning is a technique aimed at reducing the size and complexity of neural networks by eliminating redundant weight parameters (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015). Its application in LLMs has been somewhat limited, primarily due to challenges that arise in managing sparse matrices (Frantar & Alistarh, 2022; Sun et al., 2023; Frantar & Alistarh, 2023; Zhang et al., 2024). This complexity becomes particularly evident when using modern GPU hardware, as these systems are typically optimized for operations involving dense matrices (Wang, 2020; Shi et al., 2020).
@@
In the realm of LLMs, a significant similarity in output is observed among successive transformer blocks (Din et al., 2023; Liu et al., 2023). This similarity arises because each transformer block incrementally contributes to the residual path spanning the entire LLM. Figure 1 depicts the typical architecture of conventional LLMs, characterized by a continuous stack of transformer blocks. A key aspect of LLM computation is the residual path that traverses the entire network, a feature introduced to stabilize the backpropagation process during training. Consequently, each transformer block contributes its computational outputs, derived from both the attention mechanisms and feed-forward layers, to this residual path. This design feature results in a considerable degree of output similarity between consecutive transformer blocks, leading to redundancy within the LLMs.
@@ 2.1 Pruning
Standard LLMs are designed as shown in Figure 1, with a series of identically structured transformer blocks arranged sequentially. Each transformer block in these models consists of attention and feed-forward layers. The operations in both the linear and feed-forward layers primarily involve matrix multiplication between the weight parameters and input activations. Given this structure, there are two main strategies for developing compact and fast LLMs: enhancing the efficiency of each individual transformer block or reducing the overall number of blocks. To improve both the compactness and processing speed of individual blocks, pruning is frequently utilized (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015). This technique involves the strategic removal of superfluous weight parameters, thereby reducing the computational load associated with matrix multiplication in each linear and feed-forward layer. However, the application of pruning techniques to LLMs encounters a notable challenge.
Challenge 1) Limitation in Achieving LLM Inference Speedup: Pruning can be divided into two main types: unstructured and structured (Frantar & Alistarh, 2022; Sun et al., 2023; Frantar & Alistarh, 2023; Zhang et al., 2024). Unstructured pruning targets the removal of individual weights, leading to sparse weight matrices within the model. This sparsity can create complex data access patterns, complicating the management of sparse data and potentially hindering the model’s acceleration. Particularly with NVIDIA GPUs, achieving a speedup through unstructured pruning typically requires reaching a high level of sparsity, often more than 90% (Wang, 2020; Shi et al., 2020). However, for LLMs, achieving pruning ratios above 50% is often difficult without substantially compromising their linguistic prowess. Structured pruning, on the other hand, involves the elimination of predefined units of weights to create more hardware-friendly patterns, such as 2:4 pruning (Sun et al., 2023; Frantar & Alistarh, 2023; Zhang et al., 2024) or channel-wise (a.k.a. row/column-wise) pruning (Ashkboos et al., 2024; Ma et al., 2023). The goal of structured pruning is to form dense matrices that are more efficiently processed by hardware, thereby enhancing efficiency. Nonetheless, the anticipated speedup from structured pruning, ideally proportional to the pruning ratio, frequently fails to be realized in practical settings.
@@
Recent NVIDIA GPUs support the acceleration of 2:4 fine-grained structured sparsity using their Tensor Cores (Mishra et al., 2021). This form of sparsity involves having two zero values within each contiguous block of four values, resulting in a natural sparsity rate of 50%. The peak performance of sparse tensor cores is twice that of dense tensor cores. Thus, previous research on pruning LLMs has extensively examined 2:4 pruning techniques (Sun et al., 2023; Frantar & Alistarh, 2023; Zhang et al., 2024). However, it is challenging to achieve the desired speedup with 2:4 pruning. The speedup achieved with 2:4 pruning heavily depends on the size of the input matrix. Figure 3 provides a comparative analysis of matrix multiplication latency on an NVIDIA RTX A6000 GPU, using both NVIDIA cuBLAS and CUTLASS libraries. The matrix multiplication involves batched inputs with dimensions of $b\times h$, along with a weight matrix size of $h\times h$, where $b$ represents the batch size and $h$ represents the hidden size. The measurement results indicate that within a realistic batch size range of 1 to 128, 2:4 pruning does not result in a speedup. However, as the batch size is increased beyond this range, 2:4 pruning begins to show a speedup, eventually reaching approximately 1.5$\times$ the speed of dense matrix multiplication. To achieve significant speedups, it is essential to have a sufficiently large matrix size, which can lead to high arithmetic intensity (Mishra et al., 2021). In summary, achieving faster LLM inference with conventional pruning methods has two main limitations. Firstly, it requires dedicated hardware, such as the 2:4 sparse matrix multiplication accelerator adapted in recent NVIDIA GPUs, to enhance inference speed. Secondly, even with dedicated hardware support, the extent of speedup achievable through 2:4 pruning is influenced by factors like batch size and model size.
@@
As an alternative to traditional pruning methods, approaches like LLM-Pruner (Ma et al., 2023) and SliceGPT (Ashkboos et al., 2024) have been developed, focusing on the elimination of redundant neurons and the weight parameters connected to them. These methods involve the elimination of the entire channel (row/column) from weight matrices, allowing pruned weights to remain in a dense format. This retention of dense matrices enables direct speedup in matrix multiplication. However, it is challenging to maintain linguistic capabilities of LLMs with these channel-wise pruning approaches without the help of extensive fine-tuning (Ma et al., 2023). Additionally, speedup in matrix multiplication does not always lead to corresponding improvement in the end-to-end inference speed of LLMs (Ashkboos et al., 2024). This limitation arises because transformer blocks in LLMs encompass various operations beyond matrix multiplication that involve weight parameters. These operations include, but are not limited to, layer normalization and attention mechanisms (Figure 1).
Recently, Deja Vu (Liu et al., 2023) has introduced a new pruning approach termed contextual sparsity. This method dynamically assesses whether to bypass certain segments of a layer operation, based on the context of the input. It has shown effectiveness in accelerating inference, particularly in single-batch scenarios. However, its dynamic nature introduces a set of challenges akin to those encountered with early exit strategies. These challenges will be further examined and discussed in the following subsection.
@@ 2.2 Early Exit
Reducing the total number of transformer blocks in the LLM can directly lead to increased processing speed, proportional to the reduction in the number of blocks. A notable technique in this context is the early exit (Schuster et al., 2022; Din et al., 2023; Varshney et al., 2023; Chen et al., 2023). This method dynamically assesses whether to process subsequent transformer blocks. When the model reaches a certain level of confidence in its early outputs, it can halt further processing, effectively reducing computational cost. Another recent approach (Del Corro et al., 2023) proposes an alternative method for bypassing transformer blocks. This strategy suggests that removing blocks, particularly those at the beginning of LLMs, might be more feasible. However, all of these strategies, which involve skipping transformer blocks, typically require either dynamic decision-making or extensive training to be effective. To assess the impact of such removals without dynamic decision-making and training, we analyze the perplexity scores of OPT-6.7B/13B on the WikiText-2 dataset (Merity et al., 2016) after removing various consecutive blocks, as illustrated in Figure  4. The analysis results show that the “Early Exit” approach leads to a noticeable decline in the linguistic prowess of the LLMs, as evidenced by a significant increase in their perplexity scores. Moreover, “Chunk Best”, which represents the best perplexity results achieved by testing all possible removable points of consecutive blocks, also results in a significant increase in perplexity as the number of removed blocks increases. Therefore, the concept of removing consecutive blocks from LLMs is not effective without dynamic decision-making and training. In addition, applying early exit to LLMs, particularly in real-world implementation, faces the following challenges.
@@ 3.1 Output Similarity across Transformer Blocks
Here, $x_{i}$ denotes the output of the $i$-th transformer block, and $T_{i}$ denotes the operation results of the $i$-th transformer block. To comprehensively analyze and quantify the output similarity within the transformer blocks, we analyze cosine similarity between outputs of these blocks. This analysis is based on the distance measurement method outlined in (Din et al., 2023). It involves feeding a single random token into each LLM and measuring the cosine similarity between the output of the $i$-th transformer block and that of the $j$-th transformer block is defined as follows:
@@ 4.1 Experimental Setup
We implement SLEB in PyTorch (Paszke et al., 2019), using the HuggingFace Transformers library (Wolf et al., 2020). The experiments on redundancy verification and elimination of transformer blocks are executed on NVIDIA A100 GPUs equipped with 80GB of memory. SLEB requires 2 GPUs for pruning OPT-66B and LLaMA-70B, and 1 GPU for pruning smaller models. Runtime of SLEB for pruning LLMs falls within the range of runtime observed with other pruning methods (See Appendix A.1). For example, SLEB can fully compress LLaMA-2-70B, within approximately 1.5 hours. SLEB is processed based on the inference-only approach without any fine-tuning, aligning with contemporary practices in post-training pruning of LLMs. We use 128 samples randomly selected from WikiText-2 training dataset as calibration data, following the approach used in previous works (Ashkboos et al., 2024).
Our evaluation encompasses models from the OPT and LLaMA-2 families. We assess SLEB under two target sparsity levels: 10% and 20%. In case that the product of the total number of transformer blocks in a model and the target sparsity is not an integer, we round up the product to determine the number of transformer blocks to remove. The proposed SLEB is compared to previous works that are widely recognized for their potential to accelerate LLM inference. These previous works include SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2023), and DSnoT (Zhang et al., 2024), which utilize 2:4 structured pruning methods, and LLM-Pruner and SliceGPT, which employ channel-wise pruning.
@@ 4.3 Language Modeling
We evaluate the linguistic capabilities of pruned LLMs on language modeling tasks. Table 1 showcases the perplexity results on C4 validation dataset (Raffel et al., 2020). It is important to note that though the sparsity of SLEB is 10% to 20% whereas previous works show a sparsity ratio between 20% to 50%, as we will discuss in Section 4.6, SLEB outperforms in LLM inference speedup. Though Wanda and DSnoT completely fail on OPT-66B, SLEB sufficiently preserves perplexity scores for all evaluated cases. Moreover, SLEB outperforms previous pruning methods in the majority of cases by a significant margin, though it uses a coarser granularity of pruning unit. These results once again emphasize a point discussed in Section 3.1 - LLMs exhibit robustness at the level of transformer blocks. By effectively identifying and removing transformer blocks, SLEB can have a minimal impact on perplexity scores. The perplexity results on WikiText-2 test dataset can be found in Appendix B.2.
@@ 4.5 Zero-shot Tasks
Following the evaluation methods presented in previous works (Ashkboos et al., 2024), we evaluate SLEB on following zero-shot tasks: PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2019), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018). LM Evaluation Harness (Gao et al., 2023) with its default parameters is used for the evaluation.
@@ 4.7 Compatibility with Post-Training Quantization
Post-training quantization (PTQ) is a well-established technique for compressing and accelerating LLMs (Frantar et al., 2023; Lee et al., 2024; Lin et al., 2023). If PTQ can be applied to compress LLMs alongside SLEB, we can expect further improvements in memory efficiency and LLM inference speed. In Figure 10, we present a comparison of perplexity scores for LLMs that have undergone SLEB with target sparsity of 20% and those that have been further compressed with AWQ (Lin et al., 2023), a state-of-the-art PTQ algorithm that quantizes weights to 4-bit integer values. The experimental results indicate that applying PTQ to streamlined models has no discernible impact on their perplexity. More analysis results on compatibility with PTQ are provided in Appendix B.6.
@@ B.8 Fine-tuning
We investigate the effects of fine-tuning on LLaMA-2-7B after pruning with 2:4 pruning (SparseGPT), channel-wise pruning (SliceGPT), and SLEB. All pruned models are trained fore a single epoch on the Alpaca training dataset (Taori et al., 2023). For the efficient fine-tuning procedure, we employ LoRA (Hu et al., 2021) modules with a rank of 8.